# CaLiV: LiDAR-to-Vehicle Calibration of Arbitrary Sensor Setups via Object Reconstruction

### v1

@@ IV EXPERIMENT DESIGN
In this chapter, we present the experimental setup used to evaluate our proposed framework. The simulation is performed in the driving simulator CARLA [30] and consists of the digital twin of the research vehicle EDGAR [31], a generic calibration target, and a flat surface. To validate our approach, the experiments are also performed in a real-world environment with the research vehicle EDGAR. Furthermore, we describe the evaluation metrics of the S2S and S2V calibration performance.
IV-A Simulation Setup
In both the simulation and the real world, the vehicle is equipped with two LiDARs with a non-overlapping FoV. One LiDAR is mounted on the front end of the vehicle roof at the position $\begin{bmatrix}2&0&2\end{bmatrix}^{T}$ and is rotated $-10\degree$ around the y-axis. It, therefore, faces the driving direction and is tilted to the ground. The other LiDAR is mounted to the roof’s rear end at the position $\begin{bmatrix}-0.4&0&2\end{bmatrix}^{T}$ and is also rotated $-10\degree$ around the y-axis and additionally $180\degree$ around the z-axis. This means that the rear LiDAR is facing the opposite direction of travel and is also tilted towards the ground. These exact mounting positions are given for the simulation, but correspond to the expected positions of the real vehicle. The characteristics of the simulated sensors, listed in Table I, are chosen to match those of real-world sensors. The resulting LiDAR FoVs are qualitatively shown in the optimization part of Fig. 1.
@@ IV-A Simulation Setup
Accelerometer Noise [$\frac{m}{s^{2}}$] (x,y,z) 0.05
Gyroscope Noise [$\frac{rad}{s}$] (x,y,z) 0.005
GPS Noise [°$\degree$] (lat,lon) 1.5e-7
@@
Orientation Noise [$rad$] (x,y,z) 0.005
Velocity Noise [$\frac{m}{s}$] (x,y,z) 0.1
@@
The simulated vehicle is also equipped with an IMU (accelerometer and gyroscope) and GPS. All sensors and additional data are augmented by an artificially added Gaussian noise. As a calibration target we use a plastic chair. It should be noted, however, that any target can be used for these experiments. The only requirements are that the target should be detectable by the LiDAR sensor and for non-overlapping FoVs it is a prerequisite to use non-symmetric targets with respect to the observed perspectives. In the real-world experiments, an additional cubic target is placed at a position that is not part of the optimization as described in Section III to serve as a validation target (see Fig. 6). Another important consideration is the robot maneuver, since certain S2V errors are not observable for some trajectories. We use the work of [5] as a basis for these considerations, which primarily uses curves for S2V calibration. The trajectory used in the simulation is shown in Fig. 1 and was executed with an average velocity of $2\frac{m}{s}$ and a static steering angle of about $17\degree$. These values are chosen such that they fulfill maximum observability for the S2V calibration and require minimal space for the robot maneuver. The real-world setup is shown in Fig. 6 and is analogous to the simulated environment.
Figure 3: Left: Curved maneuver with the EDGAR [31] in the real-world setup. Right: A chair as a calibration target and the cubic validation target.
@@ IV-B Evaluation
In the simulation, the ground truth LiDAR transformations relative to the vehicle frame are known. These are denoted as ${}^{V}_{L_{1}}T$ and ${}^{V}_{L_{2}}T$ for LiDAR $L_{1}$ and $L_{2}$, respectively. After the optimization, estimates for corrected LiDAR transformations are computed. However, these corrections might still contain errors. Therefore, we define the LiDAR transformation estimates as ${}^{V_{1}}_{L_{1}}T$ and ${}^{V_{2}}_{L_{2}}T$ for $L_{1}$ and $L_{2}$, respectively. Using these four transformations, the S2S and S2V calibration errors can now be computed separately. The $L_{1}$ S2V errors can be computed directly by computing the residual of ${}^{V_{1}}_{L_{1}}T$ relative to ${}^{V}_{L_{1}}T$. Equivalently, the $L_{2}$ S2V errors are equal to the residual of ${}^{V_{2}}_{L_{2}}T$ relative to ${}^{V}_{L_{2}}T$. To compute the residual of one transformation matrix relative to another in this context, the homogeneous transformations are first converted into translation vectors and Euler angles. Then, the pairwise residuals are computed to distinguish between different translation and rotation errors. Thus, the residual of an estimated transformation is the vector of its translation and rotation residuals and is evaluated accordingly.
The evaluation of the S2S calibration results is less straightforward, since there are remaining S2V errors in ${}^{V_{1}}_{L_{1}}T$ and ${}^{V_{2}}_{L_{2}}T$. Let us define $L_{2}$ as the reference sensor and, therefore, define its S2V transformation as fixed. We now need to calculate ${}^{V_{1}}_{L_{1}}T^{*}$ so that ${}^{V_{1}}_{L_{1}}T^{*}$ contains no S2S calibration error relative to ${}^{V_{2}}_{L_{2}}T$. ${}^{V_{1}}_{L_{1}}T^{*}$ can be constructed as follows:
${}^{V_{1}}_{L_{1}}T^{*}={}^{V_{2}}_{L_{2}}T~{}{}^{V}_{L_{2}}T^{-1}~{}{}^{V}_{L_{1}}T$ (6)
In other words, ${}^{V_{1}}_{L_{1}}T^{*}$ describes the ground truth pose of $L_{1}$ relative to the erroneous vehicle frame, containing the S2V error of $L_{2}$. The S2S calibration error can now be defined as the residual of ${}^{V_{1}}_{L_{1}}T$ relative to ${}^{V_{1}}_{L_{1}}T^{*}$.
In our experiments, we added randomly generated errors to the ground truth transformations of $L_{1}$ and $L_{2}$ to get erroneous estimates. The translation errors for x, y and z are uniformly distributed in the range of $-0.1m$ to $0.1m$. The rotation errors for roll, pitch and yaw are also uniformly distributed, but in the range of $-3.0\degree$ to $3.0\degree$. In total, we run 100 simulations containing ground truth poses and erroneous UKF poses.
@@ V-A S2S Calibration
In this section, we evaluate our method regarding S2S calibration and compare the results for ground truth and UKF poses. In addition, we will compare our method with mlcc [23]. Due to the fact that no comparable target-based approach currently exists, we compare this targetless framework because it supports non-overlapping FoVs. For the sake of comparison, we will perform the motion proposed in Section IV-A for both algorithms. The work of [23] is evaluated with ground truth vehicle poses. The mean calibration errors of all executions are shown in Table II.
@@
CaLiV (UKF) 0.0011 0.0154 0.0200 0.0043 0.0070 0.0005
CaLiV (GT) 0.0124 0.0038 0.0088 0.0007 0.0006 0.0015
TABLE II: Mean S2S translation and rotation errors in $m$ and $rad$, respectively, using UKF and ground truth (GT) poses. mlcc comparison is performed using GT poses.
The adaptive voxelization-based work of [23] produces rather inaccurate calibration transformations in our experiments. The mean translation errors are up to several meters, while the mean rotation errors are in the range of $6\degree$ to $19\degree$. The main reason is that when exposed to erroneous LiDAR poses that are due to calibration errors, mlcc does not provide accurate results. In their work, Liu et al. state that the initial LiDAR extrinsic parameters should be pre-calibrated and accurate LiDAR poses must available, e.g. by first applying a hand-eye calibration or SLAM. From our results, we conclude that the initial S2V errors of up to $3\degree$ introduced in our experiment cannot be handled by mlcc and a more accurate prior calibration is necessary.
When evaluating the results of the proposed framework, we observe that the mean calibration errors of the execution with ground truth poses is not necessarily better than the execution with UKF poses in all dimensions. The mean x-error is $0.11cm$ for UKF poses, which is more accurate than the $1.24cm$ for ground truth poses. In addition, the mean yaw ($\psi$) error is around $0.03\degree$ for UKF poses and $0.09\degree$ for ground truth poses. However, all other calibration errors are still significantly more accurate when using ground truth poses. This behavior shows that the resulting calibration accuracy highly depends on the accuracy of the vehicle poses, however, the more accurate calibration results for the UKF poses indicate that also the registration and optimization algorithm has a high influence even in the presence of ground truth poses.
Fig. 4 shows the calibration errors’ distributions of our proposed algorithm for UKF and ground truth poses. Although there are outliers, they are relatively close to the median. Furthermore, the standard deviations are relatively low for all errors. The calibration with ground truth poses is especially reliable. The biggest difference between the results of UKF and ground truth poses is in the pitch ($\theta$) errors. Nevertheless, the maximum error of all calibrations with UKF poses is below $0.6\degree$ and the median is around $0.43\degree$.
Figure 4: The distribution of the S2S translation and rotation errors of CaLiV using UKF (blue) and ground truth (green) poses.
An important question when calculating the S2S calibration is why it is even needed, since the S2V calibration is performed for both sensors. Experimental results show that performing S2S calibration as a first step improves the overall accuracy of the subsequent S2V calibration. This pre-alignment improves the consistency between sensor measurements, resulting in more accurate S2V calibration results. To demonstrate the robustness of our approach against the extremest perspective differences, we chose a sensor setup that represents the worst-case scenario in terms of non-overlapping sensor coverage.
@@ V-B S2V Calibration
In the following, we present the S2V calibration results and compare them for ground truth and UKF poses. In addition, we distinguish between the $L_{1}$ and $L_{2}$ S2V calibration error. Since $L_{2}$ is the reference LiDAR, the $L_{1}$ S2V error is dependent on the $L_{2}$ S2V error and the S2S error. As described in Section III, the second stage of the calibration process makes use of the previously calibrated S2S transformations. For the S2V calibration, an additional constraint must be introduced: The remaining transformation to be optimized by the minimization algorithm is the S2V error ${}_{V}^{V_{2}}\bar{T}$. In order to compute the rotation errors independently from the translation errors, the S2V translation errors are fixed to zero.
Figure 5: The distribution of the S2V rotation errors of CaLiV using UKF (blue) and ground truth (green) poses. The distributions for $L_{1}$ are shown on the left, while the right plot shows the distributions of the $L_{2}$ errors.
This means, the minimization only optimizes the S2V rotation errors. Table III shows the mean $L_{1}$ and $L_{2}$ S2V roll ($\phi$), pitch ($\theta$) and yaw ($\psi$) errors after performing 100 calibrations. Again, the calibration accuracy is evaluated for both the UKF and ground truth poses. Since only the rotation errors are aimed to be solved here, we do not evaluate the translation errors. The roll and pitch errors are both less than $0.1\degree$. The mean S2V yaw errors of $L_{1}$ and $L_{2}$ for both UKF and ground truth poses are all below $0.5\degree$. Fig. 5 shows the distributions of 100 $L_{1}$ (left figure) and $L_{2}$ (right figure) S2V rotation errors in radian for both UKF and ground truth poses after the S2V calibration. The roll and pitch distributions have a low standard deviation and few outliers. All roll and pitch errors are below $0.5\degree$. Although the standard deviation of the yaw errors and the number of outliers are noticeably bigger, the mean is close to the median and the maximum errors are a little over $1\degree$.
LiDAR Poses $\delta\phi$ $\delta\theta$ $\delta\psi$
$L_{1}$ UKF 0.0042 0.0047 0.0082
GT 0.0013 0.0007 0.0084
$L_{2}$ UKF 0.0014 0.0024 0.0083
GT 0.0012 0.0002 0.0084
TABLE III: Mean S2V rotation errors in $rad$ of this work’s algorithm for the curved trajectory. The proposed method is evaluated for both ground truth and UKF poses.
@@
CaLiV (Ours) 0.0013 0.0007 0.0084
TABLE IV: Mean rotation errors in $rad$ for the curved trajectory. To compare the calibration performance only the calibration was performed using GT poses.
The S2V calibration performance of CaLiV is compared to the two open-source calibration methods Multi-LiCa [16] and SensorX2Car [18]. Both methods compute the partial LiDAR-to-Vehicle calibration using the ground plane to estimate the pitch and roll angles. The results are shown in Table IV. Note that Multi-LiCa and SensorX2Car do not provide the calibration transformation for all rotation angles, and therefore, only the roll estimates for Multi-LiCa and the roll and pitch angles for the SensorX2Car calibration algorithm can be compared. For the comparison with CaLiV, we use the ground truth poses to evaluate the calibration accuracy only. We outperform both S2V approaches using ground truth poses and, compared to SensorX2Car, we are slightly less accurate using UKF poses only in the pitch angle error.
We want to highlight the accuracy of the calibration with respect to the rotational components and also discuss the translation components. In the experiments, we observe that the translational $z$-component is not observable with our trajectory. In addition, our algorithm could not distinguish between translational $x$-error and yaw error. The rotational errors are considered to be the most important for autonomous driving tasks, since they increase with distance and therefore pose a safety concern [10]. More specifically, from a purely geometric point of view, a S2V yaw error of $1\degree$ corresponds to a lateral displacement of an object of nearly $2m$ at a distance of $100m$. Since translation errors remain constant over distance, we assume that there are no translation errors and focus on the rotation errors in the S2V optimization. At the cost of possible S2V translation errors, this assumption leads to more accurate S2V calibration results overall.
V-C Real World
@@ VI CONCLUSION & FUTURE WORK
The goal of this work was to provide a target-based framework for S2S and S2V calibration of multi-LiDAR systems with arbitrary FoV overlap. In addition, we aimed to develop an algorithm that is independent of environmental features, does not rely on external sensors, and supports arbitrary target shapes that can be registered from different perspectives. The main focus is to present the challenges of target-based S2S and S2V calibration and the methodology developed to address these challenges. CaLiV demonstrates high accuracy and robustness for both S2S and S2V calibration, and is the first target-based framework to solve both tasks simultaneously without the need for external sensing devices. We present a benchmark for both calibration tasks independently and compare our work with openly available calibration methods. CaLiV outperforms the current state of the art in both non-overlapping sensor calibration and S2V calibration. Another advantage of CaLiV is that it is particularly suitable for end-of-line calibration where space is limited. While our current method requires a curved trajectory to optimize S2S and S2V calibration, it remains to be investigated whether moving the target would yield equally accurate results. Furthermore, the possibility of using a more complex trajectory could be explored to also solve the S2V translation errors while maintaining the accuracy.

### v2

@@ IV EXPERIMENT DESIGN
In this chapter, we present the experimental setup used to evaluate our proposed framework. The simulation is performed in the driving simulator CARLA [28] and consists of the digital twin of the real-world vehicle, a generic calibration target, and a flat surface. To validate our approach, the experiments are also performed in a real-world environment. Furthermore, we describe the evaluation metrics of the S2S and S2V calibration performance.
IV-A Simulation & Real World Setup
In both the simulation and the real world, the vehicle is equipped with two LiDARs with a non-overlapping FoV. One LiDAR is mounted on the front end of the vehicle roof at the position $\begin{bmatrix}2&0&2\end{bmatrix}^{T}$ and is rotated $10\degree$ around the y-axis. It, therefore, faces the driving direction and is tilted to the ground. The other LiDAR is mounted on the roof’s rear end at the position $\begin{bmatrix}-0.4&0&2\end{bmatrix}^{T}$ and is also rotated $10\degree$ around the y-axis and additionally $180\degree$ around the z-axis. This means that the rear LiDAR is facing the opposite direction of travel and is also tilted towards the ground. These exact mounting positions are given for the simulation, but correspond to the expected positions of the real vehicle. The characteristics of the simulated sensors, listed in Table I, are chosen to match those of real-world sensors. The resulting LiDAR FoVs are qualitatively shown in the optimization part of Fig. 1.
@@ IV-A Simulation & Real World Setup
Accelerometer Noise $\frac{m}{s^{2}}$] (x,y,z) 0.05
Gyroscope Noise $\frac{rad}{s}$] (x,y,z) 0.005
GPS Noise $\degree$] (lat,lon) 1.5e-7
@@
Orientation Noise [radrad] (x,y,z) 0.005
Velocity Noise $\frac{m}{s}$] (x,y,z) 0.1
@@
The simulated vehicle is also equipped with an IMU (accelerometer and gyroscope) and GPS. All sensors and additional data are augmented by an artificially added Gaussian noise. As a calibration target we use a plastic chair. It should be noted, however, that any target can be used for these experiments. The only requirements are that the target should be detectable by the LiDAR sensor and, for non-overlapping FoVs, it is a prerequisite to use non-symmetric targets with respect to the observed perspectives. In the real-world experiments, an additional cubic target is placed at a position that is not part of the optimization as described in Section III to serve as a validation target (see Fig. 4). Another important consideration is the robot maneuver, since certain S2V errors are not observable for some trajectories. We use the work of [5] as a basis for these considerations, which primarily uses curves for S2V calibration. The trajectory used in the simulation is shown in Fig. 1 and was executed with an average velocity of $2\frac{m}{s}$ and a static steering angle of about $17\degree$. These values are chosen such that they fulfill maximum observability for the S2V calibration and require minimal space for the robot maneuver. The real-world setup is shown in Fig. 4 and is analogous to the simulated environment.
Figure 4: Left: Curved maneuver with the test vehicle in the real-world setup. Right: A chair as a calibration target and the cubic validation target.
@@ IV-B Evaluation
In the simulation, the ground truth LiDAR transformations relative to the vehicle frame are known and denoted as $^{V}_{L_{1}}T$ and $^{V}_{L_{2}}T$ for LiDAR $L_{1}$ and $L_{2}$, respectively. After the optimization, we recovered the optimized versions $_{L_{2}}^{V}T^{\star}$ of the $L_{2}$ S2V transformation and $_{L_{1}}^{L_{2}}T^{\star}$ of the S2S transformation. To obtain the optimized $L_{1}$ S2V transformation $_{L_{1}}^{V}T^{\star}$, Eq. 1 can be used. The S2V errors left after the optimization step can now simply be defined as the residuals of $_{L_{1}}^{V}T^{\star}$ and $_{L_{2}}^{V}T^{\star}$ relative to $^{V}_{L_{1}}T$ and $^{V}_{L_{2}}T$, respectively.
For the S2S evaluation, we again define $L_{2}$ as the reference sensor and, therefore, define its S2V transformation $^{V}_{L_{2}}T^{*}$ as fixed. This optimized transformation still contains a S2V error. The following expression describes the transformation of $L_{1}$ relative to the vehicle frame which contains no S2S error but still contains the S2V error of $^{V}_{L_{2}}T^{*}$:
$^{V}_{L_{2}}T^{*}~{}^{L_{2}}_{L_{1}}T$ (5)
The S2S calibration error can now be defined as the residual of the optimized $^{V}_{L_{1}}T^{*}$ (which still contains the S2S error) relative to the expression above, since both share the same S2V error. This way, the S2S error can be evaluated independently from the S2V error.
In our experiments, we added randomly generated errors to the ground truth transformations of $L_{1}$ and $L_{2}$ to obtain different initial errors on the calibration matrices. The translation errors for x, y and z are uniformly distributed in the range of 0.1m-0.1m to $0.1m$. The rotation errors for roll, pitch and yaw are also uniformly distributed, but in the range of $3.0\degree$ to $3.0\degree$. In total, we run 100 simulations containing noisy UKF poses.
@@ V-A S2S Calibration
In this section, we evaluate our method regarding S2S calibration and compare our method with the targetless open-source algorithm mlcc [23]. Due to the fact that no comparable target-based approach currently exists, we compare this targetless framework because it supports non-overlapping FoVs. For the sake of comparison, we will perform the motion proposed in Section IV-A for both algorithms. The work of [23] is evaluated with ground truth vehicle poses. The mean calibration errors of all executions are shown in Table II.
@@
CaLiV 0.0011 0.0154 0.0200 0.0043 0.0070 0.0005
TABLE II: Mean S2S translation and rotation errors in $m$ and radrad, respectively, using UKF poses. mlcc comparison is performed using ground truth poses.
The adaptive voxelization-based work of [23] produces rather inaccurate calibration transformations in our experiments. The mean translation errors are up to several meters, while the mean rotation errors are in the range of $6\degree$ to $19\degree$. While mlcc works very well when given accurate initial extrinsics and poses, it does not provide accurate results when exposed to erroneous LiDAR poses that are due to calibration errors. In their work, Liu et al. state that the initial LiDAR extrinsic parameters should be pre-calibrated and accurate LiDAR poses must available, e.g. by first applying a hand-eye calibration or SLAM. From our results, we conclude that the initial S2V errors of up to $3\degree$ introduced in our experiment cannot be handled by mlcc and a more accurate prior calibration is necessary.
Fig. 5 shows the calibration errors’ distributions of our proposed algorithm. Although there are outliers, they are relatively close to the median. Furthermore, the standard deviations are relatively low for all errors. The maximum error of all calibrations is below $0.6\degree$ and the median is around $0.43\degree$.
Figure 5: The distribution of the non-overlapping S2S translation and rotation errors of CaLiV.
An important question when calculating the S2S calibration is why it is even needed, since the S2V calibration is performed for both sensors. Experimental results show that performing S2S calibration as a first step improves the overall accuracy of the subsequent S2V calibration. This pre-alignment improves the consistency between sensor measurements, resulting in more accurate S2V calibration results. To demonstrate the robustness of our approach against extreme differences in perspective, we chose a sensor setup representing the worst-case scenario of non-overlapping sensor coverage, with two sensors facing opposite directions.
@@ V-B S2V Calibration
Correcting the S2V transformation errors depends heavily on the vehicle’s trajectory and the distance to the calibration target. With our proposed trajectory, we found that the S2V translation parameters could not be reliably estimated due to the parameters’ lack of observability. Therefore, CaLiV is designed to estimate only the rotation parameters. This topic will be discussed further in Section V-C.
In the following, we present the S2V calibration results and distinguish between the $L_{1}$ and $L_{2}$ S2V calibration error. Since $L_{2}$ is the reference LiDAR, the $L_{1}$ S2V error is dependent on the $L_{2}$ S2V error and the S2S error. As described in Section III, the second stage of the calibration process makes use of the previously calibrated S2S transformations. For the S2V calibration, an additional constraint must be introduced: To account for the observability issues and to compute the rotation errors independently, the S2V translation errors are fixed to zero.
Figure 6: The distribution of the S2V rotation errors of CaLiV. The distributions for $L_{1}$ are shown on the left, while the right plot shows the distributions of the $L_{2}$ errors.
Table III shows the mean $L_{1}$ and $L_{2}$ S2V roll ($\phi$), pitch ($\theta$) and yaw ($\psi$) errors after performing 100 calibrations. The roll and pitch errors are both less than $0.1\degree$. The mean S2V yaw errors of $L_{1}$ and $L_{2}$ are all below $0.5\degree$. Fig. 6 shows the distributions of 100 $L_{1}$ (left figure) and $L_{2}$ (right figure) S2V rotation errors in radian after the S2V calibration. The roll and pitch distributions have a low standard deviation and few outliers. All roll and pitch errors are below $0.5\degree$. Although the standard deviation of the yaw errors and the number of outliers are noticeably bigger, the mean is close to the median and the maximum errors are a little over $1\degree$.
LiDAR $\delta\phi$ $\delta\theta$ $\delta\psi$
$L_{1}$ 0.0042 0.0047 0.0082
$L_{2}$ 0.0014 0.0024 0.0083
TABLE III: Mean S2V rotation errors in radrad of this work’s algorithm for the curved trajectory. The proposed method is evaluated for UKF poses.
@@
CaLiV 0.0013 0.0007 0.0084
TABLE IV: Mean rotation errors in radrad for the curved trajectory. To compare the calibration performance only the calibration was performed using ground truth poses.
The S2V calibration performance of CaLiV is compared to the two open-source calibration methods Multi-LiCa [16] and SensorX2Car [18]. Both methods compute the partial LiDAR-to-Vehicle calibration using the ground plane to estimate the pitch and roll angles. The results are shown in Table IV. Note that Multi-LiCa and SensorX2Car do not provide the calibration transformation for all rotation angles, and therefore, only the roll estimates for Multi-LiCa and the roll and pitch angles for the SensorX2Car calibration algorithm can be compared. For the comparison with CaLiV, we use the ground truth poses to evaluate the calibration accuracy only.
V-C Observability
We want to highlight the accuracy of the calibration with respect to the rotational components, as well as discuss the translational components. In our experiments, we found that the z-translation component is not observable with a planar trajectory. However, the z-translation can easily be compensated for using ground plane information, as presented in [16]. Another approach to handling the z-translation was proposed by iCalib [29], which used a physical ramp to introduce the necessary excitation in that dimension. More notably, due to the nature of the trajectory, the algorithm cannot distinguish between translational $x$-error and yaw error. In experiments, where the translation errors were not constrained to zero, we observed a calibration error for the reference sensor of $\delta x=0.2792~m$ and $\delta\phi=0.0304~rad$. With constrained translation errors the yaw angle can reliably be estimated with the accuracy reported in Table III. This could also be solved by driving longer trajectories, as demonstrated in [5]. However, the target-based approach limits the length and topology of the trajectory. This creates a trade-off in which the method prioritizes compensating for rotational calibration errors. It should be noted that the error of the $y$-component of the translation vector is not subject to the observability issues and achieves a mean accuracy of 0.0160m0.0160~m.
The rotational errors are considered to be the most important for autonomous driving tasks, since they increase with distance and therefore pose a safety concern [10]. More specifically, from a purely geometric point of view, a S2V yaw error of $1\degree$ corresponds to a lateral displacement of an object of nearly $2m$ at a distance of $100m$. Since translation errors remain constant over distance, we assume that there are no translation errors and focus on the rotation errors in the S2V optimization. At the cost of possible S2V translation errors, this assumption leads to more accurate S2V calibration results overall.
V-D Real World
@@ VI CONCLUSION & FUTURE WORK
The goal of this work was to provide a target-based framework for S2S and S2V calibration of multi-LiDAR systems with arbitrary FoV overlap. In addition, we aimed to develop an algorithm that is independent of environmental features and does not rely on external sensors. CaLiV demonstrates high accuracy and robustness for both S2S and S2V calibration. CaLiV outperforms the current state of the art in both non-overlapping sensor calibration and S2V calibration. While our current method requires a curved trajectory to optimize S2S and S2V calibration, it remains to be investigated whether moving the target would yield equally accurate results. Furthermore, the possibility of using a more complex trajectory could be explored to also solve the S2V translation errors while maintaining the accuracy.